OpenStudy (kaliyah0524):

Algebra 1 |Absolute Value|

OpenStudy (kaliyah0524):

Solve for x: -3|x + 7| = -12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (callisto):

Any attempt?

OpenStudy (kaliyah0524):

Not sure where to start @Callisto

OpenStudy (unklerhaukus):

-3|x + 7| = -12 start by dividing both sides by -3

OpenStudy (kaliyah0524):

So its C or B... How do i solve it to get either the 11 or -11 @UnkleRhaukus

OpenStudy (unklerhaukus):

huh? what did you get when you divided?

OpenStudy (kaliyah0524):

|x+7|=4

OpenStudy (unklerhaukus):

good, so now you have two cases I. (x+7) = 4 II. -(x+7)=4 take it one at a time

OpenStudy (unklerhaukus):

I. x+7 = 4 now take away seven . . .

OpenStudy (kaliyah0524):

-3.. I got tht already

OpenStudy (unklerhaukus):

ok thats one solution now the second one II. -(x+7) = 4 = ( divide both sides by -1 ) = ( take away seven . . . )

OpenStudy (kaliyah0524):

-11. Thnks (:

OpenStudy (unklerhaukus):

-(x+7) = 4 x+7 = -4 ( dividing both sides by -1 ) x = -4-7 ( taking away seven . . . ) x = -11 √